OpenStudy (anonymous):

Which is the 13th term of the sequence: 1/81, 1/27, 1/9,1/3, 1

OpenStudy (anonymous):

=729

OpenStudy (anonymous):

ar^(n-1) =(1/81)*3^(13-1) =6561 Its 6561 :)